2. A car starts from rest and accelerates 6 m/s2 for 5 s, after which it travels with a constant velocity for 9 s. The brakes are then applied so that it decelerates at 4 m/s2 . Find the total distance traveled by car.


1
Expert's answer
2021-09-14T11:02:36-0400

Let's divide total distance "d" into three parts:

1. "d_1" with accelration "a_1 = 6m\/s^2" travelled during "t_1 = 5s"

2. "d_2" with constant speed "v_2" travelled during "t_2 = 9s"

3. "d_3" with accelration "a_3 = -4m\/s^2".


The distance "d_1" is given by the following kinematic equation


"d_1 = \dfrac{a_1t_1^2}{2}"

The speed of the car at the end of this distance (wich will be the constant speed for the second part) can be found from the definition of acceleration:


"v_2 = a_1t_1"

The distance "d_2" travelled with the constant speed is given as follows:


"d_2 = v_2t_2 = a_1t_1t_2"

The distance "d_3" can be found from another cinematic equation:


"d_3 = \dfrac{0^2-v_2^2}{2a_3} = -\dfrac{v_2^2}{2a_3} = -\dfrac{(a_1t_1t_2)^2}{2a_3}"

where represents the final speed of the car at the end of "d_3" (car stops).

Thus, the total distance is:


"d = d_1 + d_2 + d_3\\
d = \dfrac{a_1t_1^2}{2} + v_2t_2 - -\dfrac{(a_1t_1t_2)^2}{2a_3}\\
d = \dfrac{6\cdot 5^2}{2} + 6\cdot 5\cdot 9 - \dfrac{(6\cdot 5\cdot 9) ^2}{-2\cdot 4} \approx 9458m"

Answer. 9458m.